# Domain-Specific Code Language Models: Unraveling the Potential for HPC Codes and Tasks

### v1

@@ X Conclusion
That is why, in this paper, we decided to evaluate if we can build smaller models that can do similar, if not better, than existing LLMs on HPC-related programming tasks. We built MonoCoder using existing code-oriented LLM but by carefully selecting model size such that the resulting model can fit on commodity hardware on edge devices. Our experimental results demonstrate that such smaller models can be trained on HPC-specific languages and can perform similarly to existing LLMs in terms of perplexity and code completion tasks. Furthermore, after fine-tuning MonoCoder on the dataset containing OpenMP pragmas, we found that it can outperform the considerably larger GPT-3.5 model on the task of generating OpenMP pragmas. We hope that our results will help in democratizing applications of smaller LLMs to HPC programming problems.
XI Future Work
@@
In the near future, we intend to apply Tokompiler to C and C++ corpora and integrate more code representations, such as the data-flow graph (DFG) and the intermediate representation (IR) [68], to enhance model understanding as shown in closely related works  [69, 70]. Moreover, we believe that pre-training on the provably compilable subset of HPCorpus will enhance the correctness of model predictions for compilation-oriented tasks. Then, we intend to fine-tune those pre-trained models for HPC downstream tasks, such as OpenMP pragma generation [16, 15] and MPI domain decomposition distribution [16, 17]. In general, given the differences between general programming tasks and HPC-specific programming tasks, our research vision is to systematically analyze each and every element of existing LLMs (model architecture, dataset, etc.) and redesign them as needed for HPC-specific tasks.

### v3

@@ VII Conclusion & Future Work
Existing LLMs, such as GPT-3.5, are multi-lingual and are trained on languages unrelated to HPC. This phenomenon leads to huge model sizes and demands expensive compute resources to train. Thus, we decided to evaluate if we can build a smaller, domain-specific model that can perform similar, if not better, than existing LLMs on HPC-related programming tasks. Towards that end, we built MonoCoder using an existing code-oriented LLM but by carefully selecting the model size such that the resulting model can fit on a commodity hardware. Our experimental results demonstrate that MonoCoder, although orders of magnitude smaller in size than existing LLMs, performs similar to the existing LLMs in terms of language comprehension task (perplexity score), while outperforming them in code completion task (CodeBLEU score) for general-purpose and HPC-specific C and C++ programs.
@@
In the near future, we intend to integrate additional code representations, such as the data-flow graph (DFG) and the intermediate representation (IR) [48], to enhance model understanding as shown in the closely related works [49, 50]. Moreover, we believe that pre-training on a compilable subset of HPCorpus will enhance the performance of the model for compilation-oriented tasks (as partially demonstrated in [51]). Then, we intend to fine-tune those pre-trained models for HPC-specific downstream tasks, such as OpenMP pragma generation [46, 12, 13] and MPI domain decomposition distribution [13, 14, 47]. In general, our research vision is to systematically analyze each and every element of existing LLMs (model architecture, dataset, etc.) and redesign them as needed for HPC-specific tasks.